identify the type of conic sections that has the equation
4x^2+25y^2=100
And identfy its domain and range

Answer by lwsshak3(11628)

identify the type of conic sections that has the equation
4x^2+25y^2=100
And identify its domain and range
***
4x^2+25y^2=100
divide both sides by 100
x^2/25+y^2/4=1
This is an equation of an ellipse with horizontal major axis and center at the origin (0,0)
Domain: (-5,5)
Range: (-2,2)
see graph below:
y=((100-4x^2)/25)^.5